### v1

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work
Universal quantum operations and ancilla-based readout for tweezer clocks
Ran Finkelstein    Richard Bing-Shiun Tsai    Xiangkai Sun California Institute of Technology, Pasadena, CA 91125, USA    Pascal Scholl California Institute of Technology, Pasadena, CA 91125, USA
Su Direkci
California Institute of Technology, Pasadena, CA 91125, USA
   Tuvia Gefen California Institute of Technology, Pasadena, CA 91125, USA    Joonhee Choi California Institute of Technology, Pasadena, CA 91125, USA Department of Electrical Engineering, Stanford University, Stanford, CA 94305, USA    Adam L. Shaw California Institute of Technology, Pasadena, CA 91125, USA    Manuel Endres mendres@caltech.edu California Institute of Technology, Pasadena, CA 91125, USA
Enhancing the precision of measurements by harnessing entanglement is a long-sought goal in the field of quantum metrology 1, 2. Yet attaining the best sensitivity allowed by quantum theory in the presence of noise is an outstanding challenge, requiring optimal probe-state generation and readout strategies 3, 4, 5, 6, 7. Neutral atom optical clocks 8, leading systems for measuring time, have shown recent progress in terms of entanglement generation 9, 10, 11, but currently lack the control capabilities to realize such schemes. Here we show universal quantum operations and ancilla-based readout for ultranarrow optical transitions of neutral atoms. Our demonstration in a tweezer clock platform 12, 13, 14, 15, 9, 16 enables a circuit-based approach to quantum metrology with neutral atom optical clocks. To this end, we demonstrate two-qubit entangling gates with $>$99.35% fidelity via Rydberg interactions 17, 18, 15 and dynamical connectivity 19 for optical clock qubits, which we combine with local addressing 16 to implement universally programmable quantum circuits. Using this approach, we generate a near-optimal entangled probe state 1, 4, a cascade of Greenberger-Horne-Zeilinger (GHZ) states of different sizes, and perform dual-quadrature 5 GHZ readout. We also show repeated fast phase detection with non-destructive conditional reset of clock qubits and minimal dead time between repetitions by implementing ancilla-based quantum logic spectroscopy 20 (QLS) for neutral atoms. Finally, we extend this to multi-qubit parity checks and measurement-based Bell state preparation 21, 22, 23, 24. Our work lays the foundation for hybrid processor-clock devices with neutral atoms and more generally points to a future of practical applications for quantum processors linked with quantum sensors25.
Figure 1: Universal quantum operations for tweezer clocks. a, We demonstrate a universal quantum processor based on optical clock qubits. Such a processor is designed to perform any computational task and at the same time is a highly sensitive clock. To realize such a device, we demonstrate high-fidelity entangling gates between optical clock qubits and circuits with dynamical reconfiguration. We combine this with local single-qubit rotations via single-site shifts, and local mid-circuit readout. b, In our experiment, we use tweezer-trapped ${}^{88}$r atoms, where qubits are based on the ultranarrow optical clock transition ${}^{1}\text{S}_{0}\leftrightarrow^{3}\text{P}_{0}$ and entangling operations are realized through transient excitation to a Rydberg state $n^{3}\text{S}_{1}$. We further utilize coherent superpositions of quantized motional states in tweezers to shelve protected clock qubits for mid-circuit detection of ancilla qubits. c, We identify two instances where these universal quantum operations can be used in optimizing quantum metrology: entangled state preparation for near-optimal probes, and entanglement-assisted readout with ancilla qubits, realizing quantum logic spectroscopy for neutral atoms.
Control over quantum systems is evolving rapidly on many fronts. On the one hand, quantum processors – which use quantum mechanics to perform calculations – have matured into intermediate-scale systems with high-fidelity entangling operations and universal control capabilities 26, 27, 28, 29, 30, 19, 31. On the other hand, quantum sensors 25 and in particular optical atomic clocks 8 – which use quantum mechanics to measure time – have become both more accurate and precise, revealing new phenomena 32 and measuring relativistic effects at a laboratory scale 33, 34. It is thus a natural question, and a long-sought goal, how to utilize quantum information processing tools and entanglement to improve quantum sensors.
Two main strategies are typically considered in this context 1. First, entanglement introduces correlations between probes, resulting in jointly enhanced sensitivity; unfortunately, this also entails jointly enhanced vulnerability to errors and reduced dynamic range, often nullifying or strongly reducing any effective gains 3, 35, 4, 36, 37. Schemes put forward to circumvent this naïve drawback use specific, finely-tuned entangled probe-state preparation and readout schemes 1, 4, 6. While recently one such scheme has been realized with trapped ions 7, there is so far no demonstration in neutral atom optical clocks, despite recent progress in entanglement generation 9, 10, 11.
Second, entanglement with an ancillary system can be harnessed in non-destructive readout of a sensor to benefit metrology, e.g., logic spectroscopy of inaccessible probes as demonstrated with ions 20, sequential multi-round phase estimation 1, fast mid-circuit detection for multi-ensemble metrology 5, and shallow-circuit preparation of long-range entangled states 21, 22, 23, 24. Only very recently, ancilla-based readout for hyperfine qubits in neutral atoms has been shown 24, 38, but a demonstration for ultranarrow optical transitions – as needed for optical atomic clocks – is outstanding. Further, it is critical that ancilla-based readout can be performed repeatedly to fully realize its potential for metrology (and for quantum error correction 39), which is an outstanding challenge for neutral atoms in general.
To demonstrate such complex entangled probe-state generation and repeated ancilla-based readout, we realize a scalable universal quantum processor with neutral atom qubits encoded in an ultranarrow optical transition (Fig. 1a). Importantly, the same experimental system can be utilized as a neutral atom optical clock 8, which currently forms the most stable frequency reference 33, 34. Our system specifically is a tweezer optical clock 12, 13, 14, 15, 9, 16, composed of neutral strontium-88 trapped in a tweezer array 40, 41 (Fig. 1b,c), which we now equip with universal quantum computing and repeated ancilla-based readout capabilities.
In particular, we demonstrate high-fidelity entangling gates 27, 17, 18, 42 for optical clock qubits 15, achieving a controlled-Z (CZ) fidelity exceeding $0.9935(10)$ and generate entangled optical Bell pairs with a record fidelity. We further show optical coherence is preserved during qubit transport in tweezers, previously realized with hyperfine qubits 19, 43, enabling dynamical circuit reconfiguration. Combined with single-site control 16 and mid-circuit readout 44, this allows us to take a circuit-based approach to realize strategies for entangled state preparation and readout designed to improve metrology (Fig. 1c).
As part of this, we demonstrate simultaneous generation of a cascade of different-sized entangled GHZ states towards a quantum metrology scheme 1, 4 that circumvents the trade-off between entanglement-enhanced sensitivity and loss of dynamic range. This is particularly relevant in the context of state-of-the-art neutral atom optical clocks for which a loss of dynamic range implies an effective loss of stability, and where there is no expected gain, e.g., from fixed-sized GHZ states. Similarly, naive utilization of squeezed states is expected to yield only marginal gains 36, 37, 7. We also demonstrate ancilla-based mid-evolution readout, effectively realizing quantum logic spectroscopy 20 for neutral atoms, with numerous possible applications. First, this enables readout and reset with very low dead time, which could minimize the Dick effect, a current limitation of neutral atom optical clocks 8. Second, fully realizing the aforementioned quantum metrology scheme 1, 4 also requires implementing classically optimal clock interrogation 5, 4. While there is recent progress towards this end using dynamical decoupling 16, an advantageous avenue in terms of robustness to frequency noise consists of fast mid-evolution readout and reset, as demonstrated here. Third, we demonstrate ancilla-based readout for multi-qubit observables, which we use for a measurement-based generation of GHZ states, enabling in principle generation of large-scale GHZ states with shallow circuits 21, 22, 23, 24. These results more generally point to practical applications of quantum computers integrated in or connected with quantum sensors, as discussed in more detail in the outlook.
High-fidelity entangling gates for optical qubits We begin by demonstrating high-fidelity entangling gates for the ultranarrow optical clock transition. The optical transition defines a qubit composed of the ground state of strontium-88 $\mathrm{{}^{1}S_{0}}$, denoted as $|0\rangle$, and the metastable state $\mathrm{{}^{3}P_{0}}$, denoted as $|1\rangle$ (Fig. 1b). To realize high-fidelity entangling operation, we apply a single, phase-modulated, Rydberg pulse for a CZ gate, which was shown to be time-optimal 42 and recently implemented in hyperfine and nuclear atomic qubits 17, 18. To simultaneously realize high-fidelity single-qubit rotations, we operate at a magnetic field of 450 G which balances between sufficiently high Rabi frequency on the clock transition (scaling linearly with the magnetic field 45) and sufficiently strong interactions (as the Rydberg interaction strength is magnetic field dependent, see Methods, Ext. Data Fig. 1). We prepare atoms in the $|1\rangle$ state and perform erasure-cooling 44 which prepares atoms close to the motional ground state.
Figure 2: High-fidelity entangling gates for optical clock qubits. a, Global randomized benchmarking (RB) circuit with a fixed number $\mathrm{N_{rep}}$ of random single-qubit gates $R_{C}$ and a variable number $\mathrm{N_{CZ}}$ of CZ gates. A final rotation $R_{f}$ is precomputed to maximize the return probability $P_{|11\rangle}$ in the absence of errors. b, By fitting the return probability as a function of $\mathrm{N_{CZ}}$, we infer an RB fidelity of $0.9935(10)$. We find this to be consistent with a full error model which predicts an RB fidelity of $0.9943(12)$ (shaded area denotes model uncertainty), accounting for both CZ errors and single-qubit errors. Without single-qubit errors, our model predicts a CZ fidelity of $0.9963(2)$, indicating the true experimental CZ fidelity might be underestimated by $2(1)\cdot 10^{-3}$ due to residual sensitivity of the RB sequence to single-qubit errors. c, A circuit for the creation and the characterization of the optical Bell state $|\Phi^{+}\rangle$. d, Parity oscillation with a measured bare parity contrast of $0.963^{+7}_{-10}$. Together with the population overlap (Methods), we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected). e, Array configuration depicted by average atomic fluorescence image. The interatomic separation is chosen to induce strong Rydberg interactions within a pair and avoid unwanted interactions between pairs.
Figure 3: Simultaneous generation of a cascade of GHZ states and GHZ dual-quadrature readout. a, A quantum circuit for the simultaneous preparation and characterization of GHZ states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$, here $j=0,1,2$. b, Array reconfiguration steps depicted by average atomic fluorescence images at each array configuration (1), (2), (3), as marked in the circuit in a. Dashed arrows represent the atom movements between two array configurations. c, Simultaneous parity oscillations for $j=0,1,2$, from top to bottom, showing a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity. d, Simultaneous dual-quadrature parity readout of two copies of a 4-atom GHZ state realized through a $\pi/8$ phase shift on each atom in one copy, resulting in a collective $\pi/2$ phase shift before the final readout pulse.
We benchmark the performance of such gates by performing global two-qubit randomized benchmarking (RB) with single-qubit rotations 18, 17 drawn from the Pauli group. By design, the RB circuit in Fig. 2a results in maximal return probability of both atoms to the clock state ($P_{|11\rangle}$) in the case of perfect entangling gates and single-qubit rotations. Varying the number of CZ gates, $\mathrm{N_{CZ}}$, we infer an RB fidelity of $0.9935(10)$ by fitting the decay of the return probability as a function of $\mathrm{N_{CZ}}$ to an exponential trend (Fig. 2b).
The two-qubit RB circuit used here includes a fixed number of single-qubit pulses, to be minimally sensitive to single-qubit errors. However, we find that it has residual sensitivity to noise acting at the single-qubit level and thus likely underestimates the true CZ fidelity. Specifically, a detailed numerical model predicts an RB fidelity of $0.9963(2)$ including only two-qubit errors, which drops to $0.9943(12)$ if single-qubit errors are added. The model prediction, with both types of errors included (dashed line and shaded area in Fig. 2b), is consistent with the experimental RB result. Moreover, our model is independently found to be consistent with experimental results at the single-qubit level for a variety of test circuits (Methods), including single-qubit RB sequence with experimental fidelity of $0.9978(4)$ per $\pi/2$ pulse. The combination of experiment and consistent modeling results suggests that the true CZ fidelity is $2(1)\cdot 10^{-3}$ higher than $0.9935(10)$.
We note that independent of this correction, our result is only the second reported demonstration for neutral atoms generally, and the first in alkaline-earth atoms, to achieve two-qubit gate fidelity above 0.99 with Rydberg-based gates for long-lived qubits 17.
As an important application of the CZ gate, we generate an array of Bell pairs with optical qubits (Fig. 2c). After initializing atoms in superposition states and applying a CZ-gate, a $\pi/4$ pulse rotates the resulting quantum state into the even-parity, metrologically relevant, Bell state $|\Phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$. For the state obtained in the experiment, we measure the population overlap with the ideal Bell state as well as drive parity oscillations (Fig. 2d) to infer the off-diagonal (coherence) terms 46. Overall, we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected). This is consistent with a CZ fidelity of 0.996 and single-qubit fidelity of 0.998 per $\pi/2$ pulse from our model prediction.
Simultaneous GHZ cascade generation and dual-quadrature readout
Having demonstrated high-fidelity gates on optical clock qubits, we now turn to preparing larger entangled states and executing algorithms requiring more complex circuits with single-site addressing and dynamical reconfiguration. Specifically, we first demonstrate the preparation of a complex quantum register with metrological significance based on a cascade of entangled GHZ states. Such GHZ states enable, in principle, the best sensitivity allowed by quantum theory - the Heisenberg limit - with precision scaling as $1/N$, where $N$ is the atom number. However, for a single $N$-qubit GHZ state, while the sensitivity to phase variations is increased by a factor of $N$, the dynamic range is likewise reduced by a factor of $N$ 3, 4, 35. A similar trade-off exists for spin-squeezed entangled states 36, 37.
Specifically, in current neutral atom optical atomic clocks, a smaller dynamic range also implies effectively reduced clock stability. Dynamic range here refers to the maximum interval over which phase excursions of a clock laser (the local oscillator) can be measured and corrected for. Since phase excursions grow with time, a small dynamic range means a shorter allowable interrogation time. As the precision of phase estimation is reduced with shorter interrogation time, a smaller dynamic range results in a loss of precision. Therefore, in practice, the nominally improved sensitivity of GHZ states is counteracted by the loss of dynamic range, and no improvement of clock stability from fixed-$N$ GHZ states is expected over optimal interrogation with uncorrelated atoms 3, 1, 4, 35. For spin-squeezed states, similarly, only moderate improvements are predicted 36, 37, 7.
A proposed scheme 1, 4 to circumvent this trade-off between sensitivity and dynamic range requires the simultaneous preparation of a cascade of $M$ different groups of GHZ-like states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$ with $j=0,1...M-1$. Such a probe state enables Heisenberg-limited sensitivity, up to logarithmic corrections, while preserving the dynamic range. It thus enables a viable path to quantum-enhanced metrology under realistic conditions where local oscillator noise is the main limitation, as in current neutral atom optical clocks 47, 35. We note that these proposals are quantum-enhanced extensions of the classical optimal interrogation schemes 5, 48 that were recently experimentally shown 16, 49.
We realize a quantum circuit (Fig. 3a) for the required simultaneous preparation of GHZ states with $j=0,1,2$, utilizing our universal quantum processor aided by dynamical array reconfiguration (Fig. 3b). We note that this specific circuit is constrained in a one-dimensional system to produce GHZ states of up to four atoms. However, with further local operations that are available in our processor, one can in principle prepare arbitrary-size GHZ states with deeper circuits (Ext. Data Fig. 5). We also discuss a straightforward extension to two-dimensional arrays in the outlook, which would enable creating several copies of each GHZ state. We verify the simultaneous preparation of increasingly phase-sensitive states by varying the phase of a final $\pi/2$ pulse and measuring the parity signal $\mathcal{P}=\prod_{i}{Z_{i}}$ for each GHZ state. In Fig. 3c, we plot this signal against phase and find a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity, consistent with the 1:2:4 ratio expected.
In this demonstration, only one quadrature of the optical phase signal is measured. Importantly, the dynamic range can be doubled if both phase quadratures are probed simultaneously, so-called dual-quadrature readout 4, 5, which has been realized recently with uncorrelated optical clock qubits 16, 49. With extended dynamic range, the maximal interrogation time is increased, and hence the sensitivity grows. We demonstrate such dual-quadrature readout for entangled states (Fig. 3d) by preparing two 4-atom GHZ states, and then imparting a phase rotation to only one through local phase shifts. This enables readout in the $y$-basis in parallel to readout in the $x$-basis. Here, a $\pi/8$ phase shift per atom ($\pi/2$ phase on the entangled four-atom register) is realized by moving 16 all four atoms by $\lambda/16\approx 44$ nm. We note such phase shifts could alternatively be realized by moving a single atom in the GHZ state by $\lambda/4$.
An important question concerns the limitations on the maximum attainable GHZ state size in terms of achievable fidelity. We hypothesize that the phase noise of our clock laser is the dominant error source, which we test with experimental and numerical investigations. In the simultaneous GHZ state preparation scheme, we achieve a bare parity contrast of $0.52(3)$ for $N$=4 (Fig. 3c). To reduce the impact of clock laser phase noise, we prepare $N$=4 and $N$=2 GHZ states using a circuit with shorter idle time, i.e. wait time during which operations are done on other registers (by removing the single site operation $Z^{\frac{1}{4}}$ shown in Fig. 3a). With this sequence, we find an improved contrast of $0.68(3)$ and fidelity of $0.71(2)$ (Ext. Data Fig. 9). We note that the contrasts from both sequences are at or above the threshold of $0.51/\sqrt{N}=$ 0.5 required for a metrological gain 50, 2 at a fixed, short interrogation time. However, they do not exceed the threshold for the cascaded-GHZ scheme, which we find is higher (Methods).
Motivated by this, we study fidelity as a function of the circuit idle time. We measure a decay in fidelity and a modification of the bitstring distribution (Ext. Data Fig. 9) which follows the prediction of our error model, consistent with our hypothesis that clock laser phase noise is the dominant error source. We further run a model using a frequency noise spectrum of a 26-mHz clock laser system 51, for which we predict a $N$=4 GHZ state fidelity $\approx$ 0.97 (Methods, Ext. Data Fig. 6). Thus, we expect significantly improved results with an upgraded clock laser system, reaching beyond the metrological gain threshold for the cascaded-GHZ scheme, even for larger $N$ (Methods, Ext. Data Fig. 7).
Figure 4: Repeated ancilla-based quantum logic spectroscopy (QLS). a, We consider three ensembles of atoms: reference atoms (R; red), clock atoms (C; purple), and ancillas (A; blue). An entangling gate maps the state of clock atoms to the state of ancilla atoms before mid-circuit readout (MCR) of the ancilla state. b, Logic spectroscopy via ancilla atoms. A Ramsey interrogation is performed in parallel on reference atoms (R) and clock atoms (C). A clock-ancilla anti-CNOT gate maps the state onto ancilla atoms (A) for readout. c, Ancilla-based $\hat{X}$ measurement. A non-destructive measurement of the state $|\psi\rangle_{C}$ is realized by shelving clock qubits ($S$ operation) followed by ancilla MCR. We show that following unshelving ($S^{\dagger}$), the state of clock qubits is projected to a well-defined state in the measurement basis, conditioned on the ancilla state. d, $\rm{N_{cyc}}$ repeated rounds of QLS with minimal (2.9 ms) dead time, enabled by ancilla replacement via array reconfiguration. Each round $j$ consists of a fixed interrogation time $T_{j}$ followed by ancilla-based $\hat{X}$ measurement before a final round where the clock qubit is directly measured. e, We find improved contrast in the final measurement round when post-selecting on ancilla results (see text) in all previous rounds (purple markers) as compared to reference clock atoms which are protected during MCR but do not interact with ancillas (red markers).
Ancilla-based quantum logic spectroscopy
Having used entanglement to prepare metrological probe states, we now turn to use it to read out the resulting state. Specifically, we demonstrate a method based on mapping quantum information from a set of sensor qubits to another set of ancillary qubits, which are measured while the sensor qubits are quickly available for sensing once more. Such repeated, non-destructive, ancilla-based measurements, which are central to quantum error detection and correction 39, as well as schemes benefiting from such measurements, have remained largely unexplored with neutral atoms.
To efficiently realize ancilla-based readout in atomic systems, information-carrying data qubits must first be protected from global imaging light and their coherence must be preserved during mid-circuit readout (MCR) of ancilla qubits. Such protection can be realized by using dual-species arrays 52, cavity-assisted readout 53, a separate readout zone 24, or shelving to states dark to imaging light 54, 55, 56. In our system, we take the last approach and have recently realized MCR 44 by shelving optical qubits into superpositions of motional states which are long-lived and dark to imaging light.
By performing high-fidelity entangling gates between information-carrying (clock) qubits and ancilla qubits before MCR (Fig. 4a), we can realize a form of quantum logic spectroscopy (QLS) with neutral atom optical clock qubits, as originally introduced for mixed-species ion traps 20. In QLS, a non-destructive, indirect readout of an atom is realized by mapping its state onto an ancillary atom, followed by a readout of the ancilla. In our case, we use local control to place certain atoms in superposition states sensitive to the clock laser detuning $\Delta$, while ancilla atoms remain in the insensitive state $|1\rangle$. A Ramsey sequence maps $\Delta$ into population imbalance of clock atoms. A subsequent clock-ancilla anti-CNOT gate then maps this signal onto the ancilla state (Fig. 4b).
To realize a non-destructive ancilla-based $\hat{X}$ measurement, we then shelve clock qubits (using motional state shelving 44) and read out the ancilla state. To this end, we ramp up the depth of ancilla traps by a factor of 5 (leaving traps with clock atoms unchanged) such that they are non-resonant with the global motional shelving drive. Following shelving of clock qubits, we perform ancilla MCR which has two important utilities. First, reading out the ancilla state enables quantum logic spectroscopy of the signal field $\Delta$. Second, conditioned on the ancilla measurement result, the clock atom is projected to a known state in the measurement basis, which enables repeated interrogation. We verify this by measuring the clock qubit state in the $x$-basis via a $\pi/2$ pulse with a variable phase $\phi$ before detection (Fig. 4c). We find the clock qubit is projected onto the |+ket|$\rangle$| + ⟩ (|ket|$\rangle$| - ⟩) state conditioned on the ancilla measurement result $|0\rangle$ ($|1\rangle$) in the $z$-basis, with a contrast of $0.84(4)$ ($0.73(4)$). Erasure excision 57, 18 is applied to discard shots where shelving has failed (Methods).
Figure 5: Ancilla parity readout and measurement-based Bell state generation. a, An ancilla qubit (middle) is prepared in parallel with a Bell pair, which is then used in Ramsey interrogation. A parity signal is measured either directly on the Bell pair or through ancilla-based parity readout. b, Direct readout of the Bell pair parity signal (pink) is compared with a single ancilla readout (orange) yielding a weight-2 parity signal oscillating at the same frequency and phase within error bars (Methods). c, A circuit for measurement-based preparation of a long-range Bell pair, composed of 3-qubit cluster state generation followed by MCR on the middle ancilla qubit. Post-selection on the ancilla result $|0\rangle$ projects the two exterior atoms to the Bell state $|\Phi^{+}\rangle$. d, Characterization of the resulting Bell state with population overlap of $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$. The parity oscillation of the resulting state yields a contrast of $0.52(3)$, resulting in an overall raw Bell state fidelity of $0.61(2)$. Erasure excision is applied to discard shots where shelving has failed. e, A proposal for a continuous clock based on a modular architecture (see Outlook for details).
Ideally, non-destructive readout enables repetitive measurements of the same qubit or same register. This can be achieved by resetting the ancilla qubit and repeating the readout cycle. In our case, we utilize an 18 $\mu$s fast imaging method 57 which is much shorter than the qubit coherence time but results in a high probability of losing the ancilla atom during imaging. We thus realize repeated rounds of measurement and reset by replacing the ancilla atom with a reserve atom, which was shelved during ancilla MCR, using dynamical reconfiguration. This yields a dead time of only 2.9 ms dominated by shelving and unshelving operations. Alternatively, we demonstrate reusing the same ancilla atom by applying a 10-ms low-loss imaging followed by re-cooling (Methods, Ext. Data Fig. 10).
We utilize ancilla-based readout to perform fast and repeated phase detection ($\hat{X}$ measurement) cycles with clock qubit reset (Fig. 4d,e). Each round $j$ consists of a fixed evolution time $T_{j}$ followed by ancilla-based readout. Then, we perform one final sensing round with a variable time $t$, where we verify the clock qubit coherence via direct readout. In this case, we choose the evolution time $T_{j}$ such that, in the absence of a phase error, clock qubits are in the |+ket|$\rangle$| + ⟩ state. In Fig. 4e, we plot the resulting Ramsey signal for up to three consecutive measurement rounds. We use information from ancilla detection in previous rounds to remove shots where the ancilla state was toggled from $|1\rangle$ to $|0\rangle$ indicating an error has occurred. We contrast this with reference qubits that did not interact with ancillas and find an improved contrast for the ancilla-based detection.
Thus, our demonstration further shows that in the presence of noise, the coherence of the information-carrying qubit can be partially restored through ancilla-based measurement. In the metrological context, such fast detection and reset is advantageous for low dead-time clock operation, minimizing the Dick effect 8. Our scheme could also find use in optimal multi-ensemble schemes where phase information from the repeatedly measured ensemble is used to correct for phase slips in the longer evolving ensemble 5.
Multi-qubit parity checks
We now generalize our experimental demonstration of ancilla-based readout to multi-qubit measurements. In a multi-qubit measurement, an ancilla is entangled with multiple qubits and then measured. For example, a parity signal from an entangled quantum register, which has enhanced sensitivity to the signal field, can be read out via a single ancilla qubit. Additionally, the measurement projects the register qubits to a well-defined parity subspace. The latter may prove especially useful for the preparation of large-scale long-range entangled states, such as GHZ states, through ancilla measurements 21, 22, 23, 24, for which we show a proof-of-principle for optical qubits. We note that multi-qubit measurements, measuring higher weight observables, are also central to stabilizer-based quantum error correction 39.
We demonstrate the lowest-order version of such schemes, a weight-2 parity measurement, where the ancilla is entangled with two different clock qubits before measurement. To do this, we first prepare an input state composed of an ancilla atom in the clock state $|1\rangle$ and two atoms in a Bell state (Fig. 5a). We then perform a $XX$ parity measurement by entangling both clock qubits with the ancilla and measuring the ancilla’s state. In Fig. 5b, we compare the direct readout of a parity signal from the Bell pair (pink) with the ancilla readout (orange). We find that indeed the single ancilla atom observable now oscillates at a frequency of twice the single atom detuning $\Delta$. This frequency, as well as the phase of the signal, is consistent with a direct readout of the Bell pair parity (Methods). We note that the contrast of these signals is largely limited by the state preparation step, which in our one-dimensional architecture includes a state exchange operation (Ext. Data Fig. 11).
We then turn to demonstrate another relevant application of the conditional parity projection in a weight-2 ancilla MCR: long-range entanglement induced by measurement 21, 22, 23, 24. We first prepare a 3-qubit cluster state and then perform MCR of the central qubit in the $x$-basis (Fig. 5c). As $ZXZ$ is a stabilizer of the cluster state, this measurement generates a Bell pair with a well-defined parity for the exterior atoms, set by the initial state preparation. When post-selecting on the ancilla result $|0\rangle$, we find a Bell state with population $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$ and a parity oscillation contrast of $0.52(3)$, yielding an overall raw fidelity of $0.61(2)$ (Fig. 5d). If instead we post-select on the ancilla result $|1\rangle$, we find the population projection is the opposite, with $0.64(2)$ in the odd parity subspace $\{|01\rangle,|10\rangle\}$. The scheme demonstrated here can be considered a heralded preparation, which could be made deterministic by a feedback operation 21, 22, 23, 24. We note that with improvements in fidelity, and with sufficiently fast feedback, this could then be readily extensible to perform measurement-based preparation of large-scale GHZ states with only a constant depth circuit 22, 21, 23. This is in contrast to linear or log-depth standard unitary circuits 31, which could be instrumental for generating large-scale optical clock GHZ states in future devices.
Outlook
  Our results demonstrate how universal quantum processing and ancilla-based readout can be integrated with a neutral atom optical clock. Extending such schemes to larger two-dimensional tweezer clock systems 14 would enable the creation of multiple copies of GHZ states at a given size. We note in particular that the schemes demonstrated here could be easily generalized to two dimensions by performing all operations in parallel for a set of rows to generate multiple copies of GHZ states, e.g., by combining an acousto-optical deflector (AOD) with a spatial light modulator. Deviations from the magic-wavelength condition due to the multi-frequency nature of AODs 13 could be compensated via atom movements, as proposed in Ref. 16. We see this as a viable avenue towards quantum-enhanced operation of neutral atom clocks using cascades of GHZ states 1, 4 combined with classically optimal interrogation schemes and quadrature readout 5, 48, 16, 49. Importantly, mid-circuit ancilla-based detection or mid-circuit low-loss imaging (Ext. Data Fig. 10), as shown here, could be directly used to implement such classical optimal interrogation.
Upgrading the clock laser system used in this work to match the current state-of-the-art 33, 14 would enable generating significantly larger GHZ states than demonstrated here (Ext. Data Fig. 6, Methods). We then expect the maximum GHZ state size will eventually be limited by a combination of clock laser phase noise and CZ Rydberg gate fidelity. In this context, we note that the CZ gate fidelity is currently dominated by phase noise of the Rydberg laser (Ext. Data Fig. 3, Methods), which could easily be improved by implementing cavity filtering techniques 58. Further, utilizing native, beyond-two-qubit gates 27, 17 could reduce the required circuit depth. Finally, GHZ states can be prepared with constant depth circuits via measurement-based schemes 21, 22, for which we showed first steps with optical clock qubits in Fig. 5c,d.
Our results further point to a modular device with distinct quantum processing and sensing modules, realizing a vision of a quantum computer connected to a high-precision sensor (Fig. 5e). The processing module could be used to prepare entangled probe states in parallel to sensing, followed by a state exchange operation between the sensor and the processor. Subsequently, readout could be performed in the processing module in parallel to sensing, before preparing new entangled probe states, yielding very low dead time operation. The processor module could also perform non-trivial multi-qubit stabilizer measurements on the sensing qubits. Sensing and multi-qubit ancilla readout could then be cycled several times before further state preparation is needed 59. Options for fully realizing this vision could be based on a multi-zone approach with a single species 24, a dual-species 52 device composed of alkaline-earth and alkali atoms (e.g., strontium and rubidium), an optical-metastable-ground-state qubit architecture 55, 60 or combinations thereof. Similar capabilities could be implemented to improve precision measurements with molecules 61. Several of our results can be viewed as proof-of-principle demonstrations towards this modular vision, including state exchange operations (Ext. Data Fig. 11) and higher-weight ancilla-based readout.
More generally, our work highlights instances where entanglement can potentially improve metrological performance in hybrid devices with sensing and quantum processing capabilities. Such devices could be part of next-generation quantum technologies from field-transportable sensors 62 to state-of-the-art optical clocks 33 in the quest to push the limits of precision measurements.
Note — During the completion of this manuscript, we became aware of a related work 63.
Acknowledgements
We acknowledge fruitful discussions with Chi Zhang. We acknowledge support from the Army Research Office MURI program (W911NF2010136), the NSF QLCI program (2016245), from the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-1733907), the NSF CAREER award (1753386), and the DARPA ONISQ program (W911NF2010021). RF acknowledges support from the Troesh postdoctoral fellowship. RBST acknowledges support from the Taiwan-Caltech Fellowship. XS acknowledges support from the Todd Alworth Larson Fellowship. TG acknowledges support from the Quantum Science and Technology Scholarship of the Israel Council for Higher Education. JC acknowledges the support from the Terman Faculty Fellowship at Stanford.